OpenStudy (anonymous):

two solutions of an equation are.(6,5) and (8,5) what is the equation for the ordered pairs?

OpenStudy (madhu.mukherjee.946):

first find the slope y2-y1/x2-x1=5-5/8-6=0

OpenStudy (madhu.mukherjee.946):

now equation of the line- y-y1=m(x-x1) y-5=0(x-6) y-5=0

OpenStudy (anonymous):

Thnx for the help!

OpenStudy (madhu.mukherjee.946):

welcome!

OpenStudy (madhu.mukherjee.946):

where's my medal??????????